# Language Models Learn Rare Phenomena from Less Rare Phenomena: The Case of the Missing AANNs

### v1

@@ Abstract
Language models learn rare syntactic phenomena, but it has been argued that they rely on rote memorization, as opposed to grammatical generalization. Training on a corpus of human-scale in size (100M words), we iteratively trained transformer language models on systematically manipulated corpora and then evaluated their learning of a particular rare grammatical phenomenon: the English Article+Adjective+Numeral+Noun (AANN) construction (“a beautiful five days”). We first compared how well this construction was learned on the default corpus relative to a counterfactual corpus in which the AANN sentences were removed. AANNs were still learned better than systematically perturbed variants of the construction. Using additional counterfactual corpora, we suggest that this learning occurs through generalization from related constructions (e.g., “a few days”). An additional experiment showed that this learning is enhanced when there is more variability in the input. Taken together, our results provide an existence proof that models learn rare grammatical phenomena by generalization from less rare phenomena. Code available at https://github.com/kanishkamisra/aannalysis.
@@
Figure 1: A sampling of some of our key experimental manipulations and resulting performance changes. We iteratively train LMs on varied input corpora and measure the learning of the rare aann (“a beautiful five days”) or related constructions. E.g., we train on the default BabyLM corpus, a corpus in which we remove all aanns, a corpus in which we replace all aanns with a corrupted version (“beautiful a five days”) and measure learning of the corrupted version, and a corpus in which we remove aanns and also remove related constructions like “a few weeks is”.
@@ 1.1 Motivation and Prior Work
Humans come to learn and use rare grammatical structures, even if they have encountered those structures only rarely or even not at all (Pullum and Scholz, 2002; Pearl, 2022). For instance, humans accept the grammaticality of the PiPP construction (“surprising though it may be…”) even where the preposed element crosses a finite close boundary (“surprising though I know it may be that…”) (Pullum, 2017) and even though they may plausibly have never encountered such a sentence in their linguistic experience. [^1] Potts (2023) found that the finite-clause-crossing PiPP occurred in only 56 sentences out of a corpus of over 7 billion sentences, which is sufficiently rare that an adult English speaker may never have encountered this construction! How people come to know an utterance is grammatical has occupied a central place in linguistics. Specifically, mastery of never-before-encountered grammatical structures has been taken to mean that people are endowed with innate linguistic knowledge (Chomsky, 1986, 1957, 1965).
@@
Recent evidence, though, suggests that Large Language Models (LLMs) can learn complex grammar (Wilcox et al., 2018; Futrell et al., 2019; Linzen et al., 2016; Mahowald et al., ) even from human-scale amounts of input (Warstadt et al., 2023; Eldan and Li, 2023; Huebner et al., 2021). This raises the possibility that input data, along with an appropriately sophisticated or weakly biased statistical learning mechanism, is sufficient for learning rare constructions by allowing for models to emergently learn appropriate grammatical abstraction (Baroni, 2022; Piantadosi, 2023).
But modern LLMs often have access to much more training input than people do and thus might memorize rare constructions in a way that humans can’t (Warstadt and Bowman, 2022; Warstadt et al., 2023). The possibility that LLMs are “stochastic parrots” (Bender et al., 2021) that rely heavily on memorization, much more so than humans, is a common criticism of using LLMs to study human language (e.g., Chomsky et al., 2023).
@@
There are different levels of memorization, though, requiring different levels of abstraction. Consider the relatively rare AANN construction: “a beautiful five days in Texas” (Solt, 2007; Keenan, 2013; Dalrymple and King, 2019), which is interestingly distinct from the default “five beautiful days in Texas”. A model that strictly memorizes this construction might come to know that “a beautiful five days in Texas” is grammatical but have no idea that “a beautiful four days in Texas” is grammatical if it never appeared in its training. A model that generalizes just a bit more might know that “a beautiful five days in New York” is also grammatical by analogizing that a U.S. state can fill the slot. One that generalizes even more might find “a beautiful three hours at the mall” to be acceptable. Knowing that “an astonishing 200 pages” is acceptable requires generalization beyond mere lexical items. And knowing that “a blue five pencils” is not acceptable (because the adjective is “stubbornly distributive”, Schwarzschild 2011) requires yet more generalization. Even for an idealized learner, it is difficult to precisely formulate how many examples are required to get the right kinds of generalizations for even this one construction.
@@
There is increasing evidence that LLMs can generate novel linguistic utterances (McCoy et al., 2023), and also make subtle judgments on relatively rare constructions like these (Weissweiler et al., 2022; Potts, 2023), including the AANN construction (Mahowald, 2023). If they do so by memorizing examples verbatim from an unrealistically large input corpus, that would not be particularly informative for human processing. But, if they do learn rare grammatical phenomena from smaller amounts of data and can generalize beyond just those verbatim instances, that would raise the question of how they do it. For instance, in the context of the PiPP construction, Potts (2023) speculates that the comparative construction (e.g., “They are happier than we are.”) “may be the key to all of this [i.e., learning the PiPP]” because such constructions are “incredibly common” yet share abstract structure with the PiPP. If LLMs learn rare grammatical structures in part by learning and generalizing structures from much more common constructions, that would be powerful evidence for abstraction in LLMs and raise the possibility that even quite general learners could learn very rare phenomena without strong innate priors.
@@
To that end, our goal in this paper is to study a rare grammatical phenomenon in LMs that are trained on controlled input corpora and that are (a) of human realistic scale, and (b) systematically manipulated with respect to the target constructions as well as key related constructions. Our hypothesis is that generalization abilities of LMs on such rare phenomena come from abstractions and structures learned from more frequent phenomena—and that the more frequent phenomena are the “keys to all of this.”
@@
As a case study, we focus on the aforementioned aann construction, although we highlight how the methods used here could serve as a blueprint for work on other rare grammatical phenomena. Our method is to train different instantiations of a standard transformer model on the 100M-word BabyLM corpus, which we systematically manipulate—via removal and replacement—to shed light on the extent to which frequent and related phenomena encountered during training can give rise to abstractions that allow for generalizations in LMs. To test for generalization, we subjected our different LMs to a series of acceptability tests on sentences which do not appear in the training corpus and which specifically target the special properties of the aann construction. This approach of training on a systematically manipulated corpus has been fruitfully used to debias models (Lu et al., 2020; Maudslay et al., 2019), study whether LMs rely on lexical knowledge in learning syntactic rules (Wei et al., 2021), explore the effect of permuting words on pretrained models (Sinha et al., 2021), and test whether LMs can learn languages judged to be hard for humans (Kallini et al., 2024), among others. Our experiments yielded three main findings, summarized below.
@@ 1.2 Summary of findings
First, we find BabyLM-trained LMs to successfully generalize to novel instances of the aann construction. Performance unsurprisingly drops for LMs that were trained without being exposed to even a single aann during training, but perhaps surprisingly, not by all that much—they are substantially above chance. This suggests that certain items present in the training data might give rise to LMs’ non-trivial performance in judging acceptability of aanns. This finding is further strengthened by the fact that LMs trained on counterfactual variants of the aann—e.g., anan and naan, obtained by shuffling word order and are far less likely to share structure with training data items—are unable to generalize to those constructions as well as they do to aanns.
@@
Next, we investigated what might enable LMs’ learning of the aann, by further systematically manipulating their training data to hold out utterances conforming to specific linguistic and statistical phenomena. Through our manipulations, we find LMs become increasingly worse at predicting novel instances of the aann as more frequent, non-AANN-but-AANN-related phenomena are held out. For example, phenomena such as the treatment of measure noun phrases as singular (e.g., a few days is all we need)—similar to how aanns treat a plural NP as singular—end up making unseen aanns less likely by 31% on average. Importantly, these results could not be explained simply by loss of data—LMs that were trained with these phenomena left in but without an equivalently large chunk of the training data removed were almost as good as LMs that never saw aanns. This further strengthens the conclusion that the hypothesized linguistic phenomena did indeed affect generalization of the targeted construction. Notably, LMs are largely affected by these manipulations when they do not see any aanns during training, highlighting the expected non-trivial role of encountering some instances of aanns to show stronger generalization.
@@
Finally, we characterized the aforementioned interplay between the properties of the encountered aanns and the LMs generalizations on novel instances. Here we found LMs that observed aanns with more variability on the adjective, numeral, and noun slots to show better generalization than did LMs that saw more restricted-but-repeating instances of aanns. This importantly mimicked analogous findings of inductive inference in humans across linguistics (Goldberg, 1995, 2005; Suttle and Goldberg, 2011; Baayen, 2009; O’Donnell, 2015) and cognitive psychology studies (Osherson et al., 1990; Xu and Tenenbaum, 2007).
@@
Taken together, these results provide an existence proof that a weakly biased but sophisticated general-purpose statistical learner can in fact learn rare and complex linguistic phenomena, and that this can arise as a result of the presence of key related constructions during training.
@@ 2 General Methods
Here we describe methods that we use to characterize learning of the targeted rare construction, the aann. As mentioned in Section 1, we take these to be a general blueprint which can be used to study other rare constructions.
2.1 Data and Model
Corpus
Throughout the paper, we use the BabyLM-strict corpus (Warstadt et al., 2023) as our main training set. We use this primarily because of its tractable size (100M tokens), which allows us to (1) detect and control the instances of the target construction as well as related linguistic phenomena; and (2) train a whoppingly large amount of LMs in sufficient time. Since much of our experiments involve ablating constructions present in this corpus, a majority of the LMs are trained on ablated versions of BabyLM.
Construction Detection
We use regexes over a POS-tagged version of the BabyLM training data to detect aann instances. [^2] We use spaCy for getting POS-tags, see Appendix B for the exact details of the detection pipeline. We detect 2,301 aanns in the BabyLM corpus, which accounts for about 0.02% of all utterances, making construction quite rare for learners trained on this corpus.
Model
Our LMs are instances of the OPT LM (Zhang et al., 2022), which is an autoregressive transformer architecture. Our LMs have 12 layers and attention heads, use a vocabulary size of 16,384, and are trained for a maximum of 20 epochs using the transformers library (Wolf et al., 2020). The results we report for a given LM in our experiments are averaged over three different runs (with different random seeds). We list all other hyper-parameters and architectural details in Appendix A.
@@ 2.2 Acceptability data
2.3 Measurements
Scoring
We use the Syntactic Log-odds Ratio (SLOR) proposed by Pauls and Klein (2012); Lau et al. (2017), to score sentences in our tests. Given a sentence containing a prefix followed by our target construction $\mathcal{C}{}$ and an optional suffix, slor is computed as the log of the ratio between the probability of the construction given the prefix as estimated by the LM, and that estimated by a unigram model, normalized by the length of the construction. Given a language model $m$ and a unigram estimator $u$:
@@ 2.3 Measurements
Table 1: Well-formed and corrupted examples of the aann construction and its counterfactual versions (anan and naan). Corruption types are taken from Mahowald (2023). Note that in the actual data, these constructions are embedded in full sentences.
Figure 2: Accuracies on tests for aann and its counterfactuals (anan and naan), achieved by LMs trained on BabyLM with various aann-manipulations (aann as is, no aann, anan, naan). The dashed line represents chance performance (6.25%), the dot-dashed line represents accuracies for slors computed using 2- and 4-gram LMs trained on BabyLM. Accuracies for GPT-2-XL (Radford et al., 2019) and Llama-2-7B (Touvron et al., 2023) are computed using log-probabilities, since unigram frequencies were unavailable for these LMs’ corpora.
Importantly, we train the unigram estimator for a given corpus using the same tokenizer used to train our autoregressive LMs on that corpus. We use slor in lieu of the usual normalized log-probability measure to ensure that the comparison between two models cannot be explained simply by the difference in unigram frequencies brought about by our manipulations. All log-probabilities were computed using minicons (Misra, 2022).
Accuracy
Using slor values computed over test instances, we calculate the accuracy of a model as:
$\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}\left[\bigcap\limits_{k}^{4}\left\{\frac{\textsc{slor}{}_{\text{prefix}_{i}}(\mathcal{C}{}_{i})}{\textsc{slor}{}_{\text{prefix}_{i}}(\mathcal{C}{}^{k}_{i})}>1\right\}\right],$ (2)
where, $\mathbbm{1}$ is the indicator function, and $k$ enumerates all four of our corruptions, and $N$ enumerates the test instances. That is, an instance within our test set is considered to be correct iff. the score on the well-formed construction, $\mathcal{C}_{i}$, is greater than the score for all corrupted instances, $\mathcal{C}^{k}_{i}$. The accuracy, then, is the proportion of correct instances within the test set. Since this involves four pairwise comparisons, chance performance is 6.25%.

### v2

@@ Abstract
Language models learn rare syntactic phenomena, but the extent to which this is attributable to generalization vs. memorization is a major open question. To that end, we iteratively trained transformer language models on systematically manipulated corpora which were human-scale in size, and then evaluated their learning of a rare grammatical phenomenon: the English Article+Adjective+Numeral+Noun (aann) construction (“a beautiful five days”). We compared how well this construction was learned on the default corpus relative to a counterfactual corpus in which aann sentences were removed. We found that aanns were still learned better than systematically perturbed variants of the construction. Using additional counterfactual corpora, we suggest that this learning occurs through generalization from related constructions (e.g., “a few days”). An additional experiment showed that this learning is enhanced when there is more variability in the input. Taken together, our results provide an existence proof that LMs can learn rare grammatical phenomena by generalization from less rare phenomena. Data and code: https://github.com/kanishkamisra/aannalysis.
@@
Figure 1: We train LMs on varied input corpora and measure learning of the aann (“a beautiful five days”), comparing across systematically manipulated corpora. E.g. we train on a control corpus, a corpus in which we remove all aanns, a corpus in which we replace all aanns with a corrupted version (“beautiful a five days”), and a corpus in which we remove aanns and remove related constructions like “a few weeks is”. We measure learning of aanns and corrupted variants.
@@ 1.1 Motivation and Prior Work
Humans come to learn and use rare grammatical structures, even if they have encountered those structures only rarely or even not at all (Pullum and Scholz, 2002; Pearl, 2022). For instance, humans accept the grammaticality of the PiPP construction (“surprising though it may be…”) even where the preposed element crosses a finite close boundary (“surprising though I know it may be that…”) (Pullum, 2017) and even though they may plausibly have never encountered such a sentence in their linguistic experience (see Potts, 2023, for corpus estimate). How people come to know an utterance is grammatical has occupied a central place in linguistics. Specifically, mastery of never-before-encountered grammatical structures has been taken to mean that people are endowed with innate linguistic knowledge (Chomsky, 1986, 1957, 1965).
@@
Recent evidence, though, suggests that Large Language Models (LLMs) can learn complex grammar (Wilcox et al., 2018; Futrell et al., 2019; Linzen et al., 2016; Mahowald et al., 2024) even from human-scale amounts of input (Warstadt et al., 2023; Eldan and Li, 2023; Huebner et al., 2021). This raises the possibility that input data, along with an appropriately sophisticated or weakly biased statistical learning mechanism, is sufficient for learning rare constructions by allowing for models to emergently learn appropriate grammatical abstraction (Baroni, 2022; Misra and Kim, 2023). But modern LLMs often have access to much more training input than people do and thus might memorize in a way that humans cannot (Linzen, 2020; Warstadt, 2022; Warstadt et al., 2023). The possibility that LLMs are “stochastic parrots” (Bender et al., 2021), heavily reliant on memorization, is a common criticism of using LLMs to study human language (e.g., Chomsky et al., 2023).
@@
There are different levels of memorization, though, requiring different levels of abstraction. Consider the aann construction: “a beautiful five days in Texas” (Solt, 2007; Keenan, 2013; Dalrymple and King, 2019), which is rarer than the default “five beautiful days in Texas”. A model that strictly memorizes this phrase might come to know that “a beautiful five days in Texas” is grammatical but have no idea that “a beautiful four days in Texas” is grammatical if it never appeared in its training. A model that generalizes just a bit more might know that “a beautiful five days in New York” is also grammatical by generalizing that any U.S. state can fill the slot. Knowing that “an astonishing 200 pages” is acceptable requires generalization beyond mere lexical items. And knowing that “a blue five pencils” is not acceptable (because colors are “stubbornly distributive”, Schwarzschild 2011) requires yet more knowledge. Even for an idealized learner, it is difficult to precisely formulate how these kinds of generlizations emerge.
@@
There is increasing evidence that LLMs can generate novel linguistic utterances (McCoy et al., 2023), and also make subtle judgments on relatively rare constructions like these (Weissweiler et al., 2022; Potts, 2023), including the aann (Mahowald, 2023). If they do so by memorizing examples verbatim from an unrealistically large training corpus, that would not be particularly informative for human processing. But, if they do learn rare grammatical phenomena from smaller amounts of data and can generalize beyond just those verbatim instances, that would raise the question of how they do it and if it can inform theorizing about humans. For instance, in the context of the PiPP construction, Potts (2023) speculates that the comparative construction (e.g., “They are happier than we are.”) “may be the key to all of this [i.e., learning the PiPP]” because such constructions are “incredibly common” yet share abstract structure with the PiPP. If LLMs learn rare grammatical structures in part by learning and generalizing structures from much more common constructions, that would be powerful evidence for abstraction in LLMs and raise the possibility that even quite general learners could learn very rare phenomena without strong innate priors, drawing in part on the long-posited linguistic hypothesis that apparently distinct grammatical phenomena often share underlying structure.
@@
To that end, our goal in this paper is to study a relatively rare grammatical phenomenon in LMs trained on controlled input corpora that are (a) of human realistic scale, and (b) systematically manipulated with respect to the target constructions as well as key related constructions. Our hypothesis is that generalization abilities of LMs on such rare phenomena come from abstractions and structures learned from more frequent related phenomena—and that knowledge of more frequent phenomena is the “key to all of this.”
@@
As a case study, we focus on the aforementioned aann construction, although we highlight how the methods used here could serve as a blueprint for work on other phenomena. Our method is to train different instantiations of a transformer model on the 100M-word BabyLM corpus, which we systematically manipulate—via removal and replacement—to explore how frequent and related phenomena encountered during training facilitate generalization behavior in LMs. To test for generalization, we subjected our LMs to a series of acceptability tests on sentences which do not appear in the training corpus and which specifically target the special properties of the aann.
This approach of training on a systematically manipulated corpus has been used to debias models (Maudslay et al., 2019; Lu et al., 2020), explore the effect of permuting words on pretrained models (Sinha et al., 2021), and test whether LMs can learn languages judged to be hard for humans (Kallini et al., 2024). It is also becoming a fruitful method for giving causal answers to questions about syntactic learning in language models, including hypotheses about learning subject-verb agreement (Wei et al., 2021), the acquisition of negative polarity items (Jumelet et al., 2021; Weber et al., 2021), subject-auxiliary inversion (Warstadt, 2022), and the English passive alternation (Leong and Linzen, 2024). Using this “filtered pretraining” method, Patil et al. (2024) find evidence of syntactic generalization underlying models’ success on syntactic benchmarks. While this related work has largely focused on ubiquitous linguistic structures (e.g., passives, subject-verb agreement, etc.), we specifically zero in on a rare construction to explore learning in the linguistic “long tail”, where there is relatively little evidence available in the input.
@@ 1.2 Summary of findings
First, we find BabyLM-trained LMs to successfully generalize to novel instances of the aann construction. Performance unsurprisingly drops for LMs that were trained without being exposed to even a single aann during training, but perhaps surprisingly, not by all that much—they are substantially above chance. This suggests that certain items present in the training data might give rise to LMs’ non-trivial performance in judging acceptability of aanns. This finding is further strengthened by the fact that LMs trained on counterfactual variants of the aann—e.g., anan and naan, obtained by shuffling word order and are far less likely to share structure with training data items—are unable to generalize to those constructions as well as they do to aanns (one which they have not seen at all).
@@
Next, we investigated what might enable LMs’ learning of the aann, by further systematically manipulating their training data to hold out utterances conforming to specific linguistic and statistical phenomena. Through our manipulations, we find LMs become worse at predicting novel instances of the aann as more frequent, non-aann-but-aann-related phenomena are held out. For example, phenomena such as the treatment of measure noun phrases as singular (e.g., a few days is all we need)—similar to how aanns treat a plural NP as singular—end up making unseen aanns less likely by 36.5% on average. Importantly, these results could not be explained simply by loss of data—LMs that were trained with these phenomena left in but without an equivalently large chunk of the training data removed were almost as good as LMs that never saw aanns. This further strengthens the conclusion that the hypothesized linguistic phenomena did indeed affect generalization of the targeted construction. Notably, LMs are largely affected by these manipulations when they do not see any aanns during training, highlighting the expected non-trivial role of encountering some instances of aanns to show stronger generalization.
@@
Finally, we characterized the aforementioned interplay between the properties of the encountered aanns and the LMs generalizations on novel instances. Here we found LMs that observed aanns with more variability on the adjective, numeral, and noun slots to show better generalization than did LMs that saw more restricted-but-repeating instances of aanns. This importantly mimicked analogous findings of inductive inference in humans across disciplines (Osherson et al., 1990; Goldberg, 2005; Xu and Tenenbaum, 2007; Baayen, 2009; Suttle and Goldberg, 2011; O’Donnell, 2015).
@@
Taken together, these results provide an existence proof that a weakly biased but sophisticated general-purpose statistical learner can learn rare and complex linguistic phenomena, in part because of key related phenomena seen during training. While our analyses suggest potential links between “constructions” (Goldberg, 1995), our findings are also compatible with theories that think of rare phenomena as derivationally related (Chomsky, 1965) to more frequent and well-attested structures (much as Potts, 2023, posits shared syntactic structure as the key to the PiPP).
@@ 2 General Methods
2.1 Corpus
Throughout, we use the BabyLM-strict corpus (Warstadt et al., 2023) as our base training set. We use BabyLM-strict because of its human-realistic scale and tractable size (100M tokens), which allows us to (1) detect and control the instances of the target construction as well as related linguistic phenomena; and (2) train a large number of LMs in a reasonable timeframe.
2.2 Language Model
Our LMs are instances of OPT LM (Zhang et al., 2022), an autoregressive transformer architecture. Our LMs have 12 layers and attention heads, use a vocabulary size of 16,384, and are trained for a maximum of 20 epochs using the transformers library (Wolf et al., 2020). The results we report for a given LM are averaged over three different runs (with different random seeds). We list other hyper-parameters and architectural details in App. B.
2.3 Construction Detection
To detect aanns, we used a regex over a part-of-speech tagged version of BabyLM. Specifically, we started with a regex for detecting aanns and then measured its recall by hand-annotating examples (with annotations performed by the authors) found by an extremely permissive regex that looked for any “a” or “an” that appeared sequentially prior to a numeral and a plural noun in a sentence (thus likely capturing almost all aanns, albeit with very low precision). We used the hand annotations to iteratively refine our regex and handle special cases. We continued this process until, on the final set of hand annotations, we detected 17/18 instances (missing only an instance where “pound” was used instead of “pounds” due to an apparent typo—but since this violates the key plural-noun requirement of aanns, it is unclear if it counts as a genuine missed instance). Ultimately, our final regex detected 2,448 aanns in the BabyLM corpus (about 0.02% of the total 11.5M utterances). See App. C for our detailed pipeline and our recall analysis.
Even with the refined regex, we cannot guarantee perfect recall—a potential issue for claims about learning in the absence of any occurrences. To address this issue, we include controls in which we assume that we missed 300 aanns (a conservatively high number, given our recall estimate) and artificially “pollute” the data to drown out the effect of any remaining aanns. As described below, our conclusions were unchanged in this robustness analysis, suggesting our results were not driven by undetected aanns.
@@ 2.4 Acceptability data
2.5 Scoring and Accuracy
We use the Syntactic Log-odds Ratio (slor) (Pauls and Klein, 2012; Lau et al., 2017) to score sentences in our tests. Given a sentence containing a prefix followed by our target construction $\mathcal{C}{}$ and an optional suffix, slor is computed as the log of the ratio between the probability of the construction given the prefix as estimated by the LM, and that estimated by a unigram model, normalized by the length of the construction. Given a language model $m$ and a unigram estimator $u$:
@@ 2.5 Scoring and Accuracy
Table 1: Well-formed and corrupted examples of the aann construction and its counterfactual versions (anan and naan). Corruption types are taken from Mahowald (2023).
Figure 2: Accuracies on tests for aann and its counterfactuals (anan and naan), achieved by LMs trained on BabyLM with various aann-manipulations (aann as is, no aann, anan, naan).
[…]
and $\blacktriangle$ under the aann training condition are cases where training data was polluted by randomly replacing 300 aanns with anans and naans, respectively, in order to assess the impact of an imperfect aann detection system. The dashed line represents chance performance (6.25%) and the dot-dashed line represents accuracies for 2- and 4-gram LMs trained on BabyLM. Accuracies for GPT-2-XL (Radford et al., 2019) and Llama-2-7B (Touvron et al., 2023) are computed using log-probabilities, since unigram frequencies were unavailable for these LMs’ corpora.
Importantly, we train the unigram estimator for a given corpus using the same tokenizer used to train our autoregressive LMs on that corpus. We use slor in lieu of the usual normalized log-probability measure, ensuring that the comparison between two models cannot be explained simply by the difference in unigram frequencies due to our manipulations. Log-probabilities were computed using minicons (Misra, 2022). An instance within our test set is considered to be correct iff the slor values of the well-formed construction is greater than that for all four corrupted instances. The accuracy, then, is the proportion of correct instances within the test set. Since this involves four pairwise comparisons, chance performance is 6.25%.